### v1

@@ 2.1.3 Nonparametric statistical inverse problems
Nonparametric statistical inverse problems depart from the Bayesian and transport map settings by only modeling the noise $\eta$ as a random variable [83, 84, 85, 86]. The unknown parameter $u$ is thought of as some fixed, true function that generates data $y$ according to (2.1). Several examples can be formulated this way, such as derivative estimation, denoising, deconvolution, or the inverse heat equation [83, 64].
@@
where $\xi$ is a (typically Gaussian) white noise process indexed by the Hilbert space $\mathcal{Y}=L^{2}(\Omega)$. The forward operator $\mathcal{G}\colon\mathcal{U}\to\mathcal{Y}$ is usually linear in theoretical studies, although there are some exceptions [76, 87]. The parameter $\sigma$ scales the noise, and the number $N$ represents the “amount of information” available for recovery.
@@
where $\xi_{n}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{N}(0,1)$ are independent and identically distributed (i.i.d.) standard Gaussians and the $\{x_{n}\}$ are some observed design points in the domain $\Omega$. Thus, $N$ represents the sample size of available data. The finite-sample problem (2.7) is also known as statistical inverse learning [88, 89].
Theoretical analysis is easier for the continuum model (2.6). However, it must be carefully interpreted (i.e., in a weak sense) because $\lVert\xi\rVert_{\mathcal{Y}}=\infty$ almost surely if $\mathcal{Y}$ is infinite-dimensional. Thus, infinite-dimensional statistical inverse problems are characterized by “large” noise realizations. The error in reconstructions of the parameter $u$ is measured on average with respect to the noise. The parameter $u$ is assumed to belong to a bounded set of functions, typically the unit ball in some smoothness class such as a Sobolev or Besov space. To solve the inverse problem, one develops estimators that aim to minimize the worst-case average error over this set of functions. This is the minimax framework, which enables the classification of estimators as minimax optimal or not [83, Sec. 2].
@@ 2.1.4 Measure-centric formulations
There is indeed an emerging measure-centric line of research on distributional inversion [98, 99, 100, 101, 102, 103]. For example, measure-valued parameters arise in continuum formulations of optimal experimental design [104, 105, 106]. The calibration of physical models using diverse data sources leads to inverse problems where both the parameters and the measurements are measure-valued [107, 108]. Unlike Bayesian inversion from Section 2.1.1, it is worth emphasizing that in these frameworks, the underlying state space for the parameters is a space of distributions. Applying Bayesian ideas in the setting of measure-valued parameters would require a notion of posterior distribution over the probability measure space, which would be an interesting yet challenging future direction. The focus in this subsection—and other parts of the chapter more broadly—is on non-Bayesian inverse problems with observed data belonging to the space of probability measures. We now elaborate on this setting.
@@
A traditional inversion approach for (2.8) or (2.10) would involve stacking all $M$ equations on $\mathcal{Y}$ into a single equation on the product space $\mathcal{Y}^{M}$ [62]. Then, the enlarged inverse problem can be solved with Bayesian or optimization methods. However, this strategy may impose an ordering on the indices $\{1,\ldots,M\}$ that is not fundamental to the inverse problem. It also raises difficulties when applying operator learning to approximate the inverse map $Y^{M(u)}\mapsto u$ because such a map would have to be re-trained from scratch if $M=M(u)$ changes when $u$ changes. Relatedly, a naïve attempt to solve (2.9) by solving $M(\mu)$ individual inverse problems does not take into account correlations between points from the statistical model. One path forward is to work with distribution-valued data instead.
@@
The data is $\nu\in\mathscr{P}(\mathcal{Y})$ and the unknown is $\mu\in\mathscr{P}(\mathcal{U})$. There is a growing body of work on SIPs [110, 103, 111] and related methodologies, such as data-consistent inversion [112, 113, 114, 115]. These approaches extend classical inverse problem frameworks by treating both the data and the unknown parameters as probability measures and modeling their relationships through forward operators defined between metric spaces of probability spaces. Recent advances in SIPs have leveraged variational optimization in probability measure spaces, explored the effects of different statistical divergences (e.g., $f$-divergences and Wasserstein distances) used for parameter reconstruction, addressed challenges in conditional and marginal recovery, and introduced dynamical solution techniques. These methods provide robust stability estimates, convergence guarantees, regularization effects, and uncertainty quantification [116, 100, 102]. A comparison between SIPs and Bayesian inverse problems can be found in [117].
@@ 2.2 Regularization methods
More recently, operator learning, which focuses on estimating mappings between function spaces, is gaining significant attention in the context of inverse problems. Beyond the more common tasks of learning forward or inverse operators from paired data, the operator learning framework offers new opportunities to derive or incorporate regularization techniques to tackle challenging inverse problems. Section 5 will explore two key categories of methods for learning priors and regularization mechanisms: Bayesian variational inference and denoising-driven regularization, notably the plug-and-play technique.
Variational inference methods aim to approximate prior distributions by learning them from within a parametric family of probability measures. This is achieved by optimizing statistical divergence-based objective functions. In contrast, plug-and-play methods leverage the insight that denoisers inherently encode prior information. By embedding a learned denoiser into iterative optimization algorithms, typically replacing a proximal operator, plug-and-play methods regularize inverse problems without requiring an explicit prior. As we will examine, these prior-learning and regularization approaches can be interpreted within the broader framework of operator learning, where the mathematical properties of the learned operators (e.g., Lipschitz continuity, monotonicity) play a crucial role in ensuring algorithmic stability and convergence when solving inverse problems.
@@ 3 Background on operator learning
Supervised operator learning lifts finite-dimensional vector-to-vector regression to infinite-dimensional function-to-function regression. The framework designs approximation architectures at the continuum level. Consequently, these architectures exhibit robustness to the level of discretization used during numerical implementation [147, Sec. 2.3, pp. 8–9]. This is a core distinction between operator learning and traditional machine learning. Once trained, operator learning models can be deployed at different discretizations of the input and output spaces without additional effort, while at the same time producing consistent predictions. The primary motivators for operator learning are the acceleration of scientific computation and the discovery of physical laws from real data. Operator learning has been particularly successful in surrogate modeling of parametrized PDEs and their corresponding downstream tasks, such as uncertainty propagation or engineering design optimization, based on these PDE parameter-to-solution map surrogates.
Operator learning developed independently from the field of inverse problems and does not explicitly rely on ideas from that field. However, it is known that forward operator learning [148, 149, 150, 151]—and nonparametric regression more generally [152]—can be interpreted as a linear inverse problem with a non-compact forward map. The non-compact linear forward map arises from the point evaluations of the unknown and possibly nonlinear target operator at the randomly sampled input data points. This perspective has proven to be highly effective in the study of linear regression and kernel methods, where sharp error estimates are developed by combining concentration of measure techniques in Hilbert spaces with concrete manipulations of integral operators [153, 154, 155].
@@
For a more comprehensive overview of operator learning from a mathematical and theoretical perspective, we refer the reader to these recent survey articles [159, 163, 164]. The present section only covers material that is necessary for inverse problem purposes. Section 3.1 presents the basic components of an operator learning workflow, while Section 3.2 defines common neural operator architectures that appear frequently in the rest of this chapter.
@@ 3.1 Supervised learning of operators
In statistical learning, the data are assumed to be i.i.d. samples from some joint distribution over $\mathcal{X}_{1}\times\mathcal{X}_{2}$. Such an assumption may not hold for highly correlated processes, such as time series or optimal experimental designs. Nonetheless, in this section we further assume that $u_{n}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mu$ for some $\mu\in\mathscr{P}(\mathcal{X}_{1})$ and that $\xi_{n}\equiv 0$ for all $n$. Often in operator learning, the labeled data are not passively given, but instead are actively queried using the map $\mathsf{F}$ as a black box. Thus, the training data distribution $\mu$ is frequently specified by the user. While it is of considerable interest to determine optimal choices for $\mu$ [165, 11, 166, 167, 168], in practice $\mu$ is usually taken to be a Gaussian measure over $\mathcal{X}_{1}$ with a specified covariance structure such as Matérn or exponential [159, Sec. 4.1]. Then $\mu$ is sampled i.i.d. to generate the random functions $\{u_{n}\}$, and consequently $\{y_{n}\}$ after querying $u_{n}\mapsto\mathsf{F}(u_{n})$. The noise-free condition $\xi_{n}\equiv 0$ is justified when the data arises from high-fidelity numerical solvers with controlled discretization errors; this is a typical use case in operator learning.
@@
with respect to the training distribution $\mu$. Other test accuracies are defined by choosing different function spaces $X$ and loss functions $\ell$. The out-of-distribution error with respect to some $\mu^{\prime}\neq\mu$ is also relevant in applications [166, 148]; in this case, one would replace every instance of $\mu$ by $\mu^{\prime}$ in the preceding displays. See [159, Sec. 4.2.1] for other choices of loss functions or loss augmentations. The expectations in expressions such as (3.5) and (3.6) are evaluated numerically by averaging over a finite set of held-out test samples.
More broadly, operator learning as a field extends beyond the supervised setting considered here. The present chapter will touch on some unsupervised operator learning problems in Section 5. We also refer the reader to [151, Chapter 6.2] for other examples.
@@ 3.2 Common architectures
By now, there are dozens of variants of operator learning architectures. For an overview of some of these and their shared structure, see [159, 163, 164, Sec. 3] and [172, 169]. What these models share is a conceptually continuous design that decouples model parameters from any fixed discretization. Some operator learning methods of note that we do not study in detail in this chapter include continuum transformers for functions and probability measures [13, 173, 5, 174, 95, 175], function space random features [176, 169], kernel methods and Gaussian processes [171], and nonparametric linear operator parametrizations [156, 157, 165, 158, 177, 148, 149, 170]. The remainder of this subsection describes two important parametric classes: encoder-decoder networks in Section 3.2.1 and neural operators in Section 3.2.2. They correspond to a hypothesis space $\mathscr{H}$ in (3.2) given by
@@ 3.2.1 Encoder-decoder operator networks
is said to have encoder-decoder structure if the encoder $\mathscr{E}\colon\mathcal{X}_{1}\to\mathbb{R}^{d_{1}}$, approximator $\mathscr{A}\colon\mathbb{R}^{d_{1}}\to\mathbb{R}^{d_{2}}$, and decoder $\mathscr{D}\colon\mathbb{R}^{d_{2}}\to\mathcal{X}_{2}$ are all continuous maps [163, Sec. 2.4 and 4.1]. Encoder-decoder structure is a fundamental component of most universal approximation proofs for various operator learning methods [147, 178, 172, 179]. It can also be related to autoencoders on function spaces; see [163, Fig. 3] and [180, 181, 78]. Encoder-decoder networks are also the nonlinear architectures that enjoy the most comprehensive quantitative theory, including sharp approximate rates [179] and sample complexity bounds [182, 183].
The DeepONet architecture is a canonical encoder-decoder operator network [184]. Building off of earlier work in the shallow case [185], DeepONet constructs $\Psi$ in (3.8) from traditional, finite-dimensional, deep feedforward neural networks [186, Chp. 2] as follows. Let $\mathscr{E}=L\in\mathcal{L}(\mathcal{X}_{1};\mathbb{R}^{d_{1}})$ be a continuous linear map. For example, $L$ could be represented by projections onto a finite number of basis functions if $\mathcal{X}_{1}$ is a Hilbert space, general linear functionals in $\mathcal{X}_{1}^{*}$ if $\mathcal{X}_{1}$ is a Banach space, or point evaluations at $d_{1}$ fixed (or varying, see [187]) sensor points in $\Omega_{1}$ if $\mathcal{X}_{1}$ embeds into the space of continuous functions. The approximator $\mathscr{A}\colon\mathbb{R}^{d_{1}}\to\mathbb{R}^{d_{2}}$ is a deep neural network with $d_{2}$ output components $a_{j}\colon\mathbb{R}^{d_{1}}\to\mathbb{R}$ called the “branch net”. For another deep neural network with $d_{2}$ output components $\varphi_{j}\colon\Omega_{2}\to\mathbb{R}^{p_{2}}$ called the “trunk net”, the decoder is the linear map $\mathscr{D}\colon z\mapsto\sum_{j=1}^{d_{2}}z_{j}\varphi_{j}$. The $\{\varphi_{j}\}$ represent basis functions for the output space $\mathcal{X}_{2}$. Putting together the pieces, we can write the DeepONet as
@@
A related architecture, called PCA-Net [181], is based on principal component analysis (PCA) [188]. This requires $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$ to be separable Hilbert spaces, together with the conditions $\operatorname{\mathbb{E}}_{u\sim\mu}\lVert u\rVert^{2}_{\mathcal{X}_{1}}<\infty$ and $\operatorname{\mathbb{E}}_{u\sim\mu}\lVert\mathsf{F}(u)\rVert^{2}_{\mathcal{X}_{2}}<\infty$. Then $\mu$ and $\mathsf{F}_{\#}\mu$ (the pushforward of $\mu$ under $\mathsf{F}$) have well-defined covariance operators $\operatorname{Cov}(\mu)\in\mathcal{L}(\mathcal{X}_{1};\mathcal{X}_{1})$ and $\operatorname{Cov}(\mathsf{F}_{\#}\mu)\in\mathcal{L}(\mathcal{X}_{2};\mathcal{X}_{2})$ whose eigenfunctions (i.e., the PCA bases) determine the encoder $\mathscr{E}$ and decoder $\mathscr{D}$. The approximator $\mathscr{A}$ is still a neural network. Letting $\{e_{j}\}_{j\in\mathbb{N}}$ be the PCA eigenbasis of $\operatorname{Cov}(\mu)$, PCA-Net takes $\mathscr{E}=L\colon u\mapsto\{\langle e_{j},u\rangle_{\mathcal{X}_{1}}\}_{j=1}^{d_{1}}$. Letting $\{\varphi_{j}\}_{j\in\mathbb{N}}$ be the PCA eigenbasis of $\operatorname{Cov}(\mathsf{F}_{\#}\mu)$, PCA-Net is defined by

### v2

@@ 2.1.3 Nonparametric statistical inverse problems
Nonparametric statistical inverse problems depart from the nonparametric Bayesian and transport map settings by only modeling the noise $\eta$ as a random variable [87, 88, 89, 90]. Adopting a frequentist perspective, the unknown parameter $u$ is thought of as some fixed, true function that generates data $y$ according to (2.1). Several examples can be formulated this way, such as derivative estimation, denoising, deconvolution, or the inverse heat equation [87, 68].
@@
where $\xi$ is a (typically Gaussian) white noise process. The forward operator $\mathcal{G}\colon\mathcal{U}\to\mathcal{Y}$ is usually a linear map between function spaces in theoretical studies, although there are some exceptions [80, 91]. The parameter $\sigma$ scales the noise, and the number $N$ represents the “amount of information” available for recovery.
@@
where $\xi_{n}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{N}(0,1)$ are independent and identically distributed (i.i.d.) standard Gaussians and the $\{x_{n}\}$ are some observed design points in the domain of function space $\mathcal{Y}$. Thus, $N$ represents the sample size of available data. The finite-sample problem (2.8) is also known as statistical inverse learning [92, 93].
Theoretical analysis is easier for the continuum model (2.7). However, it must be carefully interpreted (i.e., in a weak sense) because $\lVert\xi\rVert_{\mathcal{Y}}=\infty$ almost surely if $\mathcal{Y}$ is infinite-dimensional; see the book [94] for more details about this technical point. Thus, infinite-dimensional statistical inverse problems are characterized by “large” noise realizations. The error in reconstructions of the parameter $u$ is measured on average with respect to the noise. The parameter $u$ is assumed to belong to a bounded set of functions, typically the unit ball in some smoothness class such as a Sobolev or Besov space. To solve the inverse problem, one develops estimators that aim to minimize the worst-case average error over this set of functions. This is the minimax framework, which enables the classification of estimators as minimax optimal or not [87, Sec. 2]. Linking back to Section 2.1.1, there is also work that performs frequentist analysis of Bayesian estimators within the minimax optimality framework [71, 72, 73, 62].
@@ 2.1.4 Measure-centric formulations
There is indeed an emerging measure-centric line of research on distributional inversion [103, 104, 105, 106, 107, 108, 109]. For example, measure-valued parameters arise in continuum formulations of optimal experimental design [110, 111, 112, 113]. The calibration of physical models using diverse data sources leads to inverse problems where both the parameters and the measurements are measure-valued [114, 115]. Unlike Bayesian inversion from Section 2.1.1, it is worth emphasizing that in these frameworks, the underlying state space for the parameters is a space of distributions. Applying Bayesian ideas in the setting of measure-valued parameters would require a notion of posterior distribution over the probability measure space, which would be an interesting yet challenging future direction. The focus in this subsection—and other parts of the chapter more broadly—is on non-Bayesian inverse problems with observed data belonging to the space of probability measures. We now elaborate on this setting.
@@
A traditional inversion approach for (2.9) or (2.11) would involve stacking all $M$ equations on $\mathcal{Y}$ into a single equation on the product space $\mathcal{Y}^{M}$ [66]. Then, the enlarged inverse problem can be solved with Bayesian or optimization methods. However, this strategy may impose an ordering on the indices $\{1,\ldots,M\}$ that is not fundamental to the inverse problem. It also raises difficulties when applying operator learning to approximate the inverse map $Y^{M(u)}\mapsto u$ because such a map would have to be re-trained from scratch if $M=M(u)$ changes when $u$ changes. Relatedly, a naïve attempt to solve the distributional inversion model (2.10) by solving $M(\mu)$ individual inverse problems does not take into account correlations between points from the statistical model. One path forward is to work with distribution-valued data instead.
@@
The formal $M=\infty$ problem implied by (2.15) is to find $\mu$ from population data $\nu=\mathsf{G}(\mu)$. However, rigorously passing to the limit $M\to\infty$ to relate the sequence of finite atom measurements and solutions to population measurements and solutions could be challenging. Such analysis would depend on assumptions placed on the statistical model (e.g., independence, misspecification) and the type of convergence desired. Taking these limits can be technical, as work in a similar setting has shown [116, 117, 118]. This is a clear direction for future work.
@@
The data is $\nu\in\mathscr{P}(\mathcal{Y})$ and the unknown is $\mu\in\mathscr{P}(\mathcal{U})$. There is a growing body of work on SIPs [120, 109, 121] and related methodologies, such as data-consistent inversion [122, 123, 124, 125]. These approaches extend classical inverse problem frameworks by treating both the data and the unknown parameters as probability measures and modeling their relationships through forward operators defined between metric spaces of probability measures. Recent advances in SIPs have leveraged variational optimization in probability measure spaces, explored the effects of different statistical divergences (e.g., $f$-divergences and Wasserstein distances) used for parameter reconstruction, addressed challenges in conditional and marginal recovery, and introduced dynamical solution techniques. These methods provide robust stability estimates, convergence guarantees, regularization effects, and uncertainty quantification [126, 106, 108]. A comparison between SIPs and Bayesian inverse problems can be found in [127].
@@ 2.2 Regularization methods
More recently, operator learning, which focuses on estimating mappings between function spaces, is gaining significant attention in the context of inverse problems. Beyond the more common tasks of learning forward or inverse operators from paired data, the operator learning framework offers new opportunities to derive or incorporate regularization techniques to tackle challenging inverse problems. Section 5 will explore two key categories of methods for learning priors and regularization mechanisms: empirical Bayes and denoising-driven regularization, notably the plug-and-play technique.
Empirical Bayes-type methods aim to approximate prior distributions by learning them from within a parametric family of probability measures. This is achieved by optimizing statistical divergence-based objective functions, which is also known as variational inference. In contrast, plug-and-play methods leverage the insight that denoisers inherently encode prior information. By embedding a learned denoiser into iterative optimization algorithms, typically replacing a proximal operator, plug-and-play methods regularize inverse problems without requiring an explicit prior. As we will examine, these prior learning and regularization approaches can be interpreted within the broader framework of operator learning, where the mathematical properties of the learned operators (e.g., Lipschitz continuity, monotonicity) play a crucial role in ensuring algorithmic stability and convergence when solving inverse problems.
@@ 3 Background on operator learning
Supervised operator learning lifts finite-dimensional vector-to-vector regression to infinite-dimensional function-to-function regression. The framework designs approximation architectures at the continuum level. Consequently, these architectures exhibit robustness to the level of discretization used during numerical implementation [157, Sec. 2.3, pp. 8–9]. This is a core distinction between operator learning and traditional machine learning. Once trained, operator learning models can be deployed at different discretizations of the input and output spaces without additional effort, while at the same time produce consistent predictions. The primary motivators for operator learning are the acceleration of scientific computation and the discovery of physical laws from real data. Operator learning has been particularly successful in surrogate modeling of parametrized PDEs and their corresponding downstream tasks, such as uncertainty propagation or engineering design optimization, based on these PDE parameter-to-solution map surrogates.
Operator learning developed independently from the field of inverse problems and does not explicitly rely on ideas from that field. However, it is known that forward operator learning [158, 159, 160, 161]—and nonparametric regression more generally [162]—can be interpreted as a linear inverse problem with a non-compact forward map. The non-compact linear forward map arises from point evaluations of the unknown and possibly nonlinear target operator at the randomly sampled input data points. This perspective has proven to be highly effective in the study of linear regression and kernel methods, where sharp error estimates are developed by combining concentration of measure techniques in Hilbert spaces with concrete manipulations of integral operators [163, 164, 165].
@@
For a more comprehensive overview of operator learning from a mathematical and theoretical perspective, we refer the reader to recent survey articles [169, 173, 174]. The present section only covers material that is necessary for inverse problem purposes. Section 3.1 presents the basic components of an operator learning workflow, while Section 3.2 defines common neural operator architectures that appear frequently in the rest of this chapter.
@@ 3.1 Supervised learning of operators
In statistical learning, the data are assumed to be i.i.d. samples from some joint distribution over $\mathcal{X}_{1}\times\mathcal{X}_{2}$. Such an assumption may not hold for highly correlated processes, such as time series or optimal experimental designs. Nonetheless, in this section we further assume that $u_{n}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mu$ for some $\mu\in\mathscr{P}(\mathcal{X}_{1})$ and that $\xi_{n}\equiv 0$ for all $n$. Often in operator learning, the labeled data are not passively given, but instead are actively queried using the map $\mathsf{F}$ as a black box. Thus, the training data distribution $\mu$ is frequently specified by the user. While it is of considerable interest to determine optimal choices for $\mu$ [175, 11, 176, 177, 178], in practice $\mu$ is usually taken to be a Gaussian measure over $\mathcal{X}_{1}$ with a specified covariance structure such as Matérn or squared exponential [169, Sec. 4.1]. Then $\mu$ is sampled i.i.d. to generate the random functions $\{u_{n}\}$, and consequently $\{y_{n}\}$ after querying $u_{n}\mapsto\mathsf{F}(u_{n})$. The noise-free condition $\xi_{n}\equiv 0$ is justified when the data arises from high-fidelity numerical solvers with controlled discretization errors; this is a typical use case in operator learning.
@@
with respect to the training distribution $\mu$. Other test accuracies are defined by choosing different function spaces $X$ and loss functions $\ell$. The out-of-distribution error with respect to some $\mu^{\prime}\neq\mu$ is also relevant in applications [176, 158]; in this case, one would replace every instance of $\mu$ by $\mu^{\prime}$ in the preceding displays. See [169, Sec. 4.2.1] for other choices of loss functions or loss augmentations. The expectations in expressions such as (3.5) and (3.6) are evaluated numerically by averaging over a finite set of held-out test samples.
More broadly, operator learning as a field extends beyond the supervised setting considered here. The present chapter will touch on some unsupervised operator learning problems in Section 5. We also refer the reader to [161, Chp. 6.2] for other examples.
@@ 3.2 Common architectures
By now, there are dozens of variants of operator learning architectures. For an overview of some of these and their shared structure, see [169, 173, 174, Sec. 3] and [182, 179]. What these models share is a conceptually continuous design that decouples model parameters from any fixed discretization. Some operator learning methods of note that we do not study in detail in this chapter include continuum transformers for functions and probability measures [13, 183, 5, 184, 100, 185], function space random features [186, 179], kernel methods and Gaussian processes [181], and nonparametric linear operator parametrizations [166, 167, 175, 168, 187, 158, 159, 180]. The remainder of this subsection describes two important parametric classes: encoder-decoder networks in Section 3.2.1 and neural operators in Section 3.2.2. They correspond to a hypothesis space $\mathscr{H}$ in (3.2) given by
@@ 3.2.1 Encoder-decoder operator networks
is said to have encoder-decoder structure if the encoder $\mathscr{E}\colon\mathcal{X}_{1}\to\mathbb{R}^{d_{1}}$, approximator $\mathscr{A}\colon\mathbb{R}^{d_{1}}\to\mathbb{R}^{d_{2}}$, and decoder $\mathscr{D}\colon\mathbb{R}^{d_{2}}\to\mathcal{X}_{2}$ are all continuous maps [173, Sec. 2.4 and 4.1]. Encoder-decoder structure is a fundamental component of most universal approximation proofs for various operator learning methods [157, 188, 182, 189]. It can also be related to autoencoders on function spaces; see [173, Fig. 3] and [190, 191, 82]. Encoder-decoder networks are also the nonlinear architectures that enjoy the most comprehensive quantitative theory, including sharp approximate rates [189] and sample complexity bounds [192, 193].
The DeepONet architecture is a canonical encoder-decoder operator network [194]. Building off of earlier work in the shallow case [195], DeepONet constructs $\Psi$ in (3.8) from traditional, finite-dimensional, deep feedforward neural networks [196, Chp. 2] as follows. Let $\mathscr{E}=L\in\mathcal{L}(\mathcal{X}_{1};\mathbb{R}^{d_{1}})$ be a continuous linear map. For example, $L$ could be represented by projections onto a finite number of basis functions if $\mathcal{X}_{1}$ is a Hilbert space, general linear functionals in $\mathcal{X}_{1}^{*}$ if $\mathcal{X}_{1}$ is a Banach space, or point evaluations at $d_{1}$ fixed (or varying, see [197, 198]) sensor points in $\Omega_{1}$ if $\mathcal{X}_{1}$ embeds into the space of continuous functions. The approximator $\mathscr{A}\colon\mathbb{R}^{d_{1}}\to\mathbb{R}^{d_{2}}$ is a deep neural network with $d_{2}$ output components $a_{j}\colon\mathbb{R}^{d_{1}}\to\mathbb{R}$ called the “branch net.” For another deep neural network with $d_{2}$ output components $\varphi_{j}\colon\Omega_{2}\to\mathbb{R}^{p_{2}}$ called the “trunk net,” the decoder is the linear map $\mathscr{D}\colon z\mapsto\sum_{j=1}^{d_{2}}z_{j}\varphi_{j}$. The $\{\varphi_{j}\}$ represent basis functions for the output space $\mathcal{X}_{2}$. Putting together the pieces, we can write the DeepONet as
@@
A related architecture, called PCA-Net [191], is based on principal component analysis (PCA) [199]. This requires $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$ to be separable Hilbert spaces, together with the conditions $\operatorname{\mathbb{E}}_{u\sim\mu}\lVert u\rVert^{2}_{\mathcal{X}_{1}}<\infty$ and $\operatorname{\mathbb{E}}_{u\sim\mu}\lVert\mathsf{F}(u)\rVert^{2}_{\mathcal{X}_{2}}<\infty$. Then $\mu$ and $\mathsf{F}_{\#}\mu$ (the pushforward of $\mu$ under $\mathsf{F}$) have well-defined trace-class covariance operators $\operatorname{Cov}(\mu)\in\mathcal{L}(\mathcal{X}_{1};\mathcal{X}_{1})$ and $\operatorname{Cov}(\mathsf{F}_{\#}\mu)\in\mathcal{L}(\mathcal{X}_{2};\mathcal{X}_{2})$ whose eigenfunctions (i.e., the PCA bases) determine the encoder $\mathscr{E}$ and decoder $\mathscr{D}$. The approximator $\mathscr{A}$ is still a neural network. Letting $\{e_{j}\}_{j\in\mathbb{N}}$ be the PCA eigenbasis of $\operatorname{Cov}(\mu)$ ordered in correspondence with its eigenvalues which we sort to be nonincreasing, PCA-Net takes $\mathscr{E}=L\colon u\mapsto\{\langle e_{j},u\rangle_{\mathcal{X}_{1}}\}_{j=1}^{d_{1}}$. Letting $\{\varphi_{j}\}_{j\in\mathbb{N}}$ be the ordered PCA eigenbasis of $\operatorname{Cov}(\mathsf{F}_{\#}\mu)$, PCA-Net is defined by
@@
DeepONet and PCA-Net have linear decoders defined through basis expansions. This severely restricts their approximation power [188, 189]. Encoder-decoder operator networks with nonlinear decoders include NOMAD [200] and Shift-DeepONet [201], which generalize DeepONet, and PARA-Net [202], which generalizes PCA-Net. NOMAD and PARA-Net both take the form
$\Psi_{\mathrm{NONLINEAR}}\bigl(u;(\theta_{\mathscr{A}},\theta_{\mathscr{D}})\bigr)(x)=Q\Bigl(\bigl\{a_{j}(Lu;\theta_{\mathscr{A}})\bigr\}_{j=1}^{d_{2}},x;\theta_{\mathscr{D}}\Bigr)\,$, (3.11)
where the neural network $Q({\,\cdot\,};\theta_{\mathscr{D}})\colon\mathbb{R}^{d_{2}}\times\Omega_{2}\to\mathbb{R}^{p_{2}}$ with parameters $\theta_{\mathscr{D}}$ represents the nonlinear decoder. Shift-DeepONet adopts the more structured parametrization
$\Psi_{\mathrm{SHIFT-DON}}(u)(x)=\sum_{j=1}^{d_{2}}a_{j}(Lu)\varphi_{j}\bigl(A_{j}(Lu)x+b_{j}(Lu)\bigr)\,.$ (3.12)
In (3.12), we suppress the dependence of the neural networks $\{a_{j},A_{j},b_{j},\varphi_{j}\}_{j=1}^{d_{2}}$ on their trainable parameters for brevity. It should be noted that the potential approximation benefits of nonlinear decoding must be balanced with the increase in computational cost [202, Sec. 3.4].